Cos25°/sin70°sin85°+cos70°/sin25°sin85°+cos85°/sin25°sin70° Evaluate.

Grade:11

1 Answers

Deepak Kumar Shringi
5 years ago
in the deno write
sin70=cos20 and sin85=cos5
now cos25=cos(20+5)=cos20cos5-sin20sin5
as the first term will become =1-tan5tan20 convert others as well in this way